Determine the amount of tax liability in the following situations. In all cases, the taxpayer is using the filing status of married filing jointly. Use the appropriate Tax Tables or Tax Rate Schedules.
A. Taxable income of $72,367 that includes a qualified dividend of $890.
B. Taxable income of $21,619 that includes a dividend of $362.
C. Taxable income of $147,300 that includes a qualified dividend of $4,410. (Round your immediate computations to 2 decimal places and final answer to the nearest dollar.)
D. Taxable income of $43,681 that includes a qualified dividend of $1,269.
E. Taxable income of $295,881 that includes a qualified dividend of $16,250. (Round your immediate computations to 2 decimal places and final answer to the nearest dollar.)
(Use the Tax Tables for taxpayers with taxable income under $100,000 and the Tax Rate Schedules for those with taxable income above $100,000.)
Answer #1

solution :-

Here,we need to determine the amount of the given

(A). Taxable income of $72,367 that includes a qualified dividend of $890:-

given data ,

Income = $72,367

qualified dividend = $890

Taxable income = 72,367 - 890

Taxable income = $71,477

tax on $ 71477 = 1845+(71477-18450)*15%

= 1845+(53027)*15%

= 1845+7954.05

=$9799.05

  

tax on $ 71477 = $9799.05

B. Taxable income of $21,619 that includes a dividend of $362:

income=$ 21619

dividend=$362

taxable income = 21619-362

= $21257

taxable income =$ 21257

C. Taxable income of $147,300 that includes a qualified dividend of $4,410:

income = $147,300

qualified dividend = $4410

taxable income = 147300 - 4410

= $142890

tax liabilities = 1845+((74900-18450)*15% )+((142890-74900)*25%)

= 1845+8467.5+16997.5

=$27310

  tax liabilities   =$27310

D. Taxable income of $43,681 that includes a qualified dividend of $1,269.

income = $43681

qulified dividend= $1269

taxable income= $43681-$1269

=$42412

tax liability = 1845+(42412-18450)*15%

= 1845+ 3594.3

=$ 5439.3

tax liability    =$ 5439.3

E. Taxable income of $295,881 that includes a qualified dividend of $16,250

income = $295881

qualifiefd dividend = $16250

taxable income = 295881-16250

= $279631

\therefore taxable income = $279631

tax liabilities = 1845+((74900-18450)*15%)+((151200-74900)*25%)+((230450-151200)*28%) + ((279631-230450)33%)

= 1845+8467.5+19075+22190+16229.73

= $67807.23

   tax liabilities   = $67807.23

note:

here , some values are taken by tax table for finding the tax liabilities.

please thum up thank you